# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v2

@@ 6.1. Experimental Setup
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019) and a very recent benchmark TableBench (Wu et al., 2024). The statistics of the datasets are shown in Table 1.
(1) WikiTQ contains complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
(2) TabFact provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Ye et al., 2023) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
(3) TabBench provides an evaluation of table question answering capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation. We evaluate AutoPrep on TabBench-FC and TabBench-NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question. The answers in TabBench could be a string or list of items.
Baselines. We consider the following two categories of baselines:
@@
NL2SQL (Rajkumar et al., 2022) first translates the question into a SQL program and then executes it to get the final answer from the table for the question. We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
@@
(2) TQA Methods with Data Prep. We also investigate four SOTA TQA methods that consider data prep tasks in their question answering process.
ICL-Prep. uses few-shot in-context learning (ICL) demonstrations to guide LLMs in generating programs for data prep, as shown in Figure 3. Subsequently, we employ NL2SQL to extract answers from the cleaned tables.
@@
Chain-of-Table (CoTable) (Wang et al., 2024b) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM. These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM. We implement CoTable using the original code from (cod, 2024a).
@@
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the golden answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer, such as pairs like 7.45 and 7.4499, 4 years and 4. Afterwards, we adopt accuracy as our evaluation metric.
Backbone LLMs. We evaluate our method using representative LLMs as backbones. For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613). For open-source LLMs, we choose Llama3 (Dubey et al., 2024) (Llama-3.1-70B-Instruct) and QWen2.5 (Yang et al., 2024) (QWen2.5-72B-Instruct) for evaluation.
Experiment Settings. We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit. For a fair comparison, we set the maximum token input of all methods as 8192. Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
@@ 6.2. Improvement of Data Prep for TQA
Exp-1: Impact of question-aware data prep on TQA performance. We integrate AutoPrep into our four TQA baselines w/o data prep, in order to investigate the impact of question-aware data prep on TQA performance. The results are reported in Table 2.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, NL2SQL shows the most substantial gains, achieving an average accuracy improvement of 12.22 on WikiTQ and 13.23 on TabFact across all LLM backbones. Similarly, NL2Py also shows notable improvements in its performance, after being integrated with AutoPrep. This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, End2End and CoT methods also show considerable performance gains. For example, End2End shows an average accuracy gain of 5.27 and 1.35 on WikiTQ and TabFact respectively. These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
Finding 1: The integration of AutoPrep improves the accuracy of TQA methods w/o data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, respectively.
Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@ 6.3. Data Prep Method Comparison
CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ $\underline{80.20}$ $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
@@
Exp-2: Comparison of AutoPrep with previous SOTA TQA methods with data prep. We compare AutoPrep with TQA methods that consider data prep tasks in their question answering process, using a single LLM agent. The results are reported in Table 3.
The results clearly demonstrate that AutoPrep achieves the new SOTA performance on both the WikiTQ and TabFact datasets across all LLM backbones. In particular, AutoPrep outperforms ICL-Prep with a significant accuracy improvement by 11.00 and 8.33 on WikiTQ and TabFact on average, respectively. Other baselines, except Dater, generally outperform ICL-Prep by introducing specific technical optimizations to data prep. For instance, ReAcTable applies the ReAct paradigm to iteratively generate data prep operations rather than producing all code at once, and it explicitly considers data prep tasks (e.g., filtering and augmentation). Despite this, AutoPrep still outperforms the second best method by considerable improvement. Specifically, when compared with CoTable, AutoPrep achieves better performance, with an average accuracy improvement by 3.05 and 1.96 on WikiTQ and TabFact, respectively.
These improvements are primarily attributed to our multi-agent framework, which effectively tackles the question-aware data prep challenge. The experimental results indicate that data prep is inherently complex and is hard to be resolved by a single, one-size-fits-all solution; instead, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly. Moreover, AutoPrep covers a broader range of data prep operations than these state-of-the-art TQA methods, thereby addressing gaps (e.g., normalization tasks) not fully explored by previous solutions.
Finding 2: AutoPrep outperforms the the previous SOTA TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact, respectively,
In addition, we find that DeepSeek achieves the better overall performance at lower API costs. Thus, we select DeepSeek as our default backbone LLM for subsequent experiments.
Exp-3: Evaluation of AutoPrep on tables with various sizes. To further investigate AutoPrep, we analyze its performance across tables of varying sizes. While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables. The results are reported in Figure 7.
(a) Small
(b) Medium
(c) Large
Figure 7. Comparison of AutoPrep and other data prep methods on tables with different sizes on the WikiTQ dataset.
@@
Finding 3: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
6.4. Evaluation on Generalization
@@
Exp-4: Evaluating generalization on unseen datasets. As discussed previously, we evaluate the generalization capabilities of AutoPrep on two TabBench datasets. Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
@@
As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods. Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capability of our method. The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
Finding 4: Utilizing tool-augmented execution enhances the generalization capabilities of AutoPrep on unseen datasets.
(a) Numerical Reasoning
(b) Fact Checking
@@ 6.5. Ablation Studies
Exp-5: Evaluation on the Planner agent. We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 4(a).
We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively. This indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using Direct Prompting.
Finding 5: Our proposed Chain-of-Clauses method is more effective in suggesting high-level operations.
(a) Case one
(b) Case two
Exp-6: Evaluation on the Programmer Agents. We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
As illustrated in Table 4(b), our proposed tool-augmented method achieves better performance compared with the code generation method. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. We also record the error ratio of these two methods, as shown in Table 4(b). The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
Finding 6: Our tool-augmented method performs better in both accuracy and cost, with fewer program errors.
@@
Exp-7: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.

### v3

@@ 6.1. Experimental Setup
Datasets. We evaluate AutoPrep using the following well-adopted datasets. The statistics of the datasets are shown in Table 1.
(1) WikiTQ (Pasupat and Liang, 2015) contains complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size.
(2) TabFact (Chen et al., 2019) provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Ye et al., 2023) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
(3) TabBench (Wu et al., 2024) provides an evaluation of TQA capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation. We evaluate the generalization of AutoPrep on FC and NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question. Specifically, we directly use the prompts in WikiTQ to evaluate the generalization capabilities of AutoPrep on TabBench.
Baselines. We consider the following three categories of baselines:
@@
NL2SQL (Rajkumar et al., 2022) first translates the question into an SQL program and then executes it to get the final answer from the table for the question. We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
@@
(2) Data Prep Baselines. We consider the following data prep baselines. For each baseline, we first perform data prep and then use the above NL2SQL to extract answers from the prepared tables.
Offline DataPrep (Off-Prep) performs offline data prep operations for all tables in TQA. To this end, we have surveyed and consolidated the offline data prep operations, such as data cleaning, value normalization and column renaming, adopted by current SOTA TQA methods (Yao et al., 2022; Cheng et al., 2023; Khot et al., 2022; Wang et al., 2024; Zhu et al., 2024) to construct a comprehensive offline data prep pipeline, by utilizing the Pandas library (McKinney, 2010) and the popular DataPrep toolkit (Peng et al., 2021b).
ICL-Prep uses few-shot ICL demonstrations to guide LLMs in generating Python programs for data prep, as shown in Figure 3.
(3) SOTA TQA Methods with Data Prep. We investigate four SOTA TQA methods considering data prep tasks for TQA task.
@@
AutoTQA (Zhu et al., 2024) uses a multi-agent framework for TQA. Since the official code for AutoTQA is not publicly available, we reimplement it under the guidance of the authors (see our repository (https://github.com/fmh1art/AutoPrep/src/model/autotqa). Given the high time and API costs, we evaluate AutoTQA on a sampled subset of 500 instances from each of WikiTQ and TabFact.
@@
Chain-of-Table (CoTable) (Wang et al., 2024) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM. These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM. We implement CoTable using the original code from (cod, 2024a).
@@
Evaluation Metrics. We consider both accuracy and cost.
Accuracy. We adopt the evaluator from Binder (Cheng et al., 2023) to address cases where program executions are semantically correct but do not exactly match the golden answers.
Cost. We measure both the time and API cost for all methods. For time cost, we ensure a stable network environment and record the end-to-end processing time for each method on a single TQA instance. For API cost, we follow the official pricing guidelines (https://api-docs.deepseek.com/quick_start/pricing/) and calculate the cost based on the default LLM backbones used.
Backbone LLMs. We evaluate our methods using representative LLMs as backbones. For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613). For open-source LLMs, we choose Llama3 (Dubey et al., 2024) (Llama-3.1-70B-Instruct) and QWen2.5 (Yang et al., 2024) (QWen2.5-72B-Instruct) for evaluation.
Experiment Settings. We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit. Moreover, for fair comparison, we set the maximum token input of all methods as 8192. Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
@@ 6.2. Improvement of Data Prep for TQA
Exp-1: Impact of question-aware data prep on TQA performance. We integrate AutoPrep into our four TQA baselines w/o data prep, and report the results in Table 2.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, NL2SQL shows the most substantial gains, achieving an average accuracy improvement of 12.22 on WikiTQ and 13.23 on TabFact across all LLM backbones. Similarly, NL2Py also shows notable improvements in its performance, after being integrated with AutoPrep. This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Derive and Normalize can solve these cases and improve the overall results.
Moreover, End2End and CoT methods also show considerable performance gains. These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables. Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@ 6.3. Data Prep Method Comparison
CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ 80.2080.20{80.20}80.20 $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
@@
Exp-2: Comparison of AutoPrep with data prep baselines. We compare AutoPrep against two representative baselines: a traditional offline data prep method Off-Prep and an LLM-based in-context learning data prep method ICL-Prep.
As shown in Table 3, AutoPrep consistently outperforms both baselines on WikiTQ and TabFact. Specifically, AutoPrep improves over Off-Prep by 10.02 and 5.55 on average, respectively. The performance gap can be attributed to the fact that Off-Prep lacks question guidance, making it ineffective in addressing question-specific issues such as missing semantics and irrelevant columns. Moreover, predefined normalization routines struggle with table heterogeneity, whereas AutoPrep can generate specialized code for such cases via LLMs. Similarly, AutoPrep outperforms ICL-Prep by 11.00 and 8.33 on average on the two datasets. This highlights that a multi-agent framework is essential for producing comprehensive data prep requirements and precise, executable programs.
Exp-3: Comparison of AutoPrep with previous SOTA TQA methods with data prep. We compare AutoPrep with TQA methods that integrate data prep tasks in their question-answering process. The results, shown in Table 3, highlight that AutoPrep sets a new SOTA performance on both the WikiTQ and TabFact datasets across LLM backbones. Although CoTable achieves the best overall performance among existing SOTA methods, AutoPrep outperforms it with an impressive average improvement of 3.05 on WikiTQ and 1.96 on TabFact. These improvements are largely attributed to our multi-agent framework, which effectively addresses the question-aware data prep challenges. Furthermore, when compared to AutoTQA, AutoPrep shows significant gains of 5.81 on WikiTQ and 2.77 on TabFact. This is because AutoTQA lacks comprehensive data prep, leading to execution errors or incorrect answers.
The results demonstrate that data prep is inherently complex and cannot be solved with a one-size-fits-all solution. Instead, a more effective strategy involves specialized LLM-based agents for each type of data prep task, coordinated by a centralized planning agent. Moreover, AutoPrep covers a broader range of data prep operations, filling gaps (e.g., Derive and Normalize) that previous solutions have not fully addressed.
@@
Moreover, we also evaluate AutoPrep on tables with various sizes. We find that, by employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model. More details on the experimental results and analysis can be found in our technical report (tec, 2025).
As DeepSeek achieves the best accuracy at lower cost, we select it as our default backbone LLM for subsequent experiments.
6.4. Evaluation on Efficiency
Figure 7. Efficiency evaluation. The $x$-axis and $y$-axis represent the time cost (in seconds) and monetary cost (in dollars) for processing a single instance, respectively. The color intensity of each scatter point reflects the overall accuracy.
Exp-4: How efficient is AutoPrep compared with other methods? To evaluate the efficiency of AutoPrep, we compare it with other TQA methods in terms of time cost, monetary cost, and accuracy. The results for both small and large tables are shown in Figure 7. As shown, AutoPrep achieves the best overall performance with insignificant time and monetary costs. Specifically, its cost is lower than all methods except Off-Prep and ICL-Prep. However, AutoPrep significantly outperforms these two baselines in accuracy, achieving improvements of at least 9.36 and 5.77 points on small and large tables, respectively. Furthermore, as table size increases, AutoPrep maintains a much more stable time and cost overhead compared to other state-of-the-art TQA methods.
This efficiency stems from the program-based data prep mechanism of AutoPrep, which avoids high latency and API expenses associated with directly processing tables via LLMs. For instance, in the column derivation task on Table $T_{1}$ (Figure 2(a)), CoTable invokes LLMs to generate a full list of country codes. In contrast, AutoPrep only generates a physical operator $\mathtt{extract}$ and executes it to produce the column, which results in significantly lower time and monetary costs, particularly on large tables.
6.5. In-Depth Exploration of System Capabilities
@@
Exp-5: Evaluating generalization on unseen datasets. As discussed in Section 6.1, we evaluate the generalization capabilities of AutoPrep on the TabBench datasets. Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
@@
As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods. Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capabilities of our method. The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
Table 4. Evaluating Extensibility for New Operation Types.
Method TransTQ WikiTQ
 origin  after
ICL-Prep 58.92 56.54 55.00
ReAcTable 51.04 64.13 63.03
AutoTQA 59.75 60.80 61.20
AutoPrep 68.88 66.09 66.28
Exp-6: Evaluating extensibility on datasets requiring new logical operation types. To further evaluate the extensibility of AutoPrep, we evaluate it on a more complex dataset involving logical operations not originally supported. Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset—TransTQ with the assistance of LLMs, which is available on our GitHub repository. Specifically, we prompt an LLM to identify tables from WikiTQ where inverse transformation operations (e.g., pivot and stack (Li et al., 2023)) can be applied. These operations are then executed to generate non-relational table formats, introducing new challenges for TQA on table transformation.
As shown in Table 4, without explicit guidance for handling table transformations, all baseline TQA methods exhibit poor accuracy. In contrast, after incorporating a new logical operation Transform, AutoPrep achieves the best performance, outperforming the second-best method by 9.31 points. Moreover, we also validate that adding this new operation does not degrade the performance of AutoPrep on the original WikiTQ dataset, confirming that its multi-agent architecture is well extensible to support the seamless integration of new data prep functionalities.
@@ 6.6. Ablation Studies
Exp-7: Evaluation on the Planner agent. We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 5(a).
We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively. This indicates the superiority of our proposed method in generating more accurate logical operations. Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the logical operation suggestion problem more challenging to be solved using Direct Prompting.
Exp-8: Evaluation on the Programmer Agents. We evaluate the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
As illustrated in Table 5(b), our proposed tool-augmented method achieves better performance compared with the code generation method. Considering the logical operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. We also record the error ratio of these two methods, as shown in Table 5(b). The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
@@
Exp-9: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Derive and Normalize and compare the performance with AutoPrep.